I am stuck on this portion of the equation:
12(x-4)-(x+2)=(x+2)(x-4)
I know that you have to take everything on the right side and put it to the left. I have tried different ways and I am not getting the answer.
The answer is x^2-13x+42
If you can show me each step to come to that answer so I can use it for future equations to go off of. Thank You

Answer by Alan3354(69443)


Quadratic_Equations/455850 (2011-05-31 12:24:32): I am stuck on this portion of the equation:
12(x-4)-(x+2)=(x+2)(x-4)
-----------------
12x - 48 - x - 2 = x^2 - 2x - 8
11x - 50 = x^2 - 2x - 8
x^2 -13x + 42 = 0
(x - 6)*(x - 7) = 0
x = 6, x = 7